OpenStudy (anonymous):

7x-2(x+1) = 6x+14 7x-2x+1=6x+14 5x+1=6x+14 5x-6x=14-1 -1x=13 divide both by -1 x= -13 negative 13 isnt an answer choice . confused ._.

OpenStudy (anonymous):

i have no idea where i messed up ._.

OpenStudy (austinl):

7x-2(x+1) = 6x+14 7x - 2x - 2 = 6x + 14 5x - 2 = 6x + 14 -x - 2 = 14 -x = 16 x = -16

OpenStudy (austinl):

You need to distribute all the way through the parenthesis. a(b + c) = ab + ac

OpenStudy (anonymous):

DANG IT . thats like .. something l always forget to do , thanks xD i feel like an idiot now .

OpenStudy (austinl):

No worries!